# Embodied Agents Meet Personalization: Exploring Memory Utilization for Personalized Assistance

### v1

@@ Abstract
Embodied agents empowered by large language models (LLMs) have shown strong performance in household object rearrangement tasks. However, these tasks primarily focus on single-turn interactions with simplified instructions, which do not truly reflect the challenges of providing meaningful assistance to users. To provide personalized assistance, embodied agents must understand the unique semantics that users assign to the physical world (e.g., favorite cup, breakfast routine) by leveraging prior interaction history to interpret dynamic, real-world instructions. Yet, the effectiveness of embodied agents in utilizing memory for personalized assistance remains largely underexplored. To address this gap, we present Memento, a personalized embodied agent evaluation framework designed to comprehensively assess memory utilization capabilities to provide personalized assistance. Our framework consists of a two-stage memory evaluation process design that enables quantifying the impact of memory utilization on task performance. This process enables the evaluation of agents’ understanding of personalized knowledge in object rearrangement tasks by focusing on its role in goal interpretation: (1) the ability to identify target objects based on personal meaning (object semantics), and (2) the ability to infer object–location configurations from consistent user patterns, such as routines (user patterns). Our experiments across various LLMs reveal significant limitations in memory utilization, with even frontier models like GPT-4o experiencing a 30.5% performance drop when required to reference multiple memories, particularly in tasks involving user patterns. These findings, along with our detailed analyses and case studies, provide valuable insights for future research in developing more effective personalized embodied agents. Project website: https://connoriginal.github.io/MEMENTO
@@ 1 Introduction
Embodied agents empowered by large language models (LLMs) have recently demonstrated remarkable success in executing object rearrangement tasks in household environments [18, 49, 44, 7, 28]. As the primary objective of embodied agents is to provide assistance to users while interacting with the physical world, leveraging LLMs’ natural language understanding and reasoning capabilities lead embodied agents to effectively interpret user instructions into sets of target object-location pairs that the agent should rearrange to successfully accomplish the task.
@@
But do such tasks truly reflect the challenges in providing meaningful assistance to the users? As illustrated in Figure 1, conventional embodied tasks predominantly focus on single-turn interactions with static and simplified instructions that the agents could simply follow without implicit reasoning to comprehend user intentions [2, 17, 44, 50]. However, for personalized embodied agents, it is important to understand personalized knowledge that users assign unique semantics to the physical world (e.g., favorite cup, breakfast routine) to interpret dynamic instructions. To provide personalized assistance, agents must effectively leverage memories that retain personalized knowledge from previous interactions—especially episodic memory, which enables the recall of specific events grounded in time and space [20]. Without such memory utilization, embodied agents require users to repeatedly provide detailed instructions, which may hinder user engagement and prevent natural human-agent interaction. Despite its importance, the effectiveness of embodied agents in utilizing episodic memory containing personalized knowledge remains largely underexplored.
In this work, we present Memento, a personalized embodied agent evaluation framework designed for comprehensive assessment of memory utilization for providing personalized assistance. To enable a thorough analysis of memory utilization, we divide the memory evaluation process into two stages. In the Memory Acquisition Stage, agents perform tasks with instructions containing personalized knowledge while accumulating the interaction history. Subsequently, the Memory Utilization Stage challenges agents to complete the same tasks as in the memory acquisition stage but with modified instructions that is difficult to succeed without referencing the previously acquired personalized knowledge. This design allows us to systematically quantify the impact of memory utilization on task performance. Building upon this evaluation process, we aim to analyze agents’ ability to understand personalized knowledge in object rearrangement by focusing on its role in goal interpretation: (1) the ability to identify target objects based on personal meaning (object semantics), and (2) the ability to infer object–location configurations from consistent user patterns, such as routines (user patterns).
@@
Figure 1: Comparison between traditional embodied tasks and personalized assistance tasks. Previous works focus on strictly following simple instructions, while personalized assistance agents must know user-specific knowledge, which require grounding in past interactions. This highlights the challenge of going beyond instruction-following toward context-aware personalized embodied agent.
Based on Memento, we evaluate embodied agents powered by a range of LLMs with varying capabilities, covering both open-source and proprietary models. Our findings reveal that even frontier LLMs struggle to utilize episodic memory with personalized knowledge, with GPT-4o exhibiting a 30.5% performance drop when required to reference multiple memories. Further analysis shows that this performance degradation is particularly pronounced in tasks involving user patterns, and that the agents are highly susceptible to irrelevant memories acting as distractors. We also conduct error and success case analyses to understand how embodied agents reference memories during task execution, providing valuable insights to guide future research in developing personalized embodied agents.
To summarize, our contributions are as follows:
  • We propose Memento, a novel personalized embodied agent evaluation framework designed to assess agents’ ability to utilize episodic memory for providing personalized assistance in object rearrangement tasks.
  • To quantify the impact of memory and analyze understanding of personalized knowledge independently of reasoning capabilities, we decompose memory usage into two stages: Memory Acquisition and Memory Utilization.
  • Through extensive experiments and analysis, we identify key limitations of current LLM-powered embodied agents in leveraging personalized knowledge from memory, and offer insights to guide future research on personalized embodied agents.
@@ 2 Related Work
LLMs have significantly advanced embodied agents’ reasoning and planning capabilities in recent years. Researchers have explored LLMs for interpreting user goals [2], high-level task planning [17], and integrating LLMs into comprehensive embodied agent frameworks [18, 29, 34, 44, 19]. Other research directions have focused on generating executable code for embodied tasks directly from language instructions [30, 47, 43], while various benchmarks have been developed to evaluate embodied reasoning abilities [26, 9, 7, 28]. Collectively, these studies highlight the promise of LLM-powered agents in bridging language understanding and physical interaction.
Memory systems for embodied agents.
@@
Previous studies on memory systems for embodied agents have primarily focused on semantic memory (e.g., scene graph, semantic map), which store and provide state information about the current environment [39, 23, 16, 51, 48], or on procedural memory (e.g., skill library) that stores action primitives, focusing on how to perform tasks to enhance the efficiency on generating the low-level action code [47, 42, 59]. Another important category is episodic memory, which captures specific past interactions and experiences with users. However, prior uses of episodic memory have mostly treated it as passive task buffers [2, 43] or histories for in-context [18, 31, 44, 8], without explicitly evaluating its role in personalized task grounding or systematic memory utilization.
Personalization for embodied agents.
@@
The importance of personalization in robotics has long been recognized [13, 25, 10], particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users. Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement [22, 49], table settings [37], or personalized object navigations [11, 4]. Recently, Xu et al. [52] aim to infer user preferences from a few demonstrations and adapt planning behavior accordingly. However, these approaches primarily focus on implicit preference adaptation or short-term reactive behaviors, without modeling user-specific knowledge in a structured manner.
3 Preliminaries
@@
We formulate the object rearrangement task for LLM-powered embodied agents as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(S,A,T,R,\Omega,O,\gamma)$, where $S$ denotes the set of environment states, $A$ is the set of actions, $T$ is the transition function, $R$ is the reward function, $\Omega$ is the observation space, $O$ is the observation function, and $\gamma$ the discount factor. At each timestep $t$, the environment is in a state $s_{t}\in S$, and the agent receives a partial observation $w_{t}\in\Omega$ in text modality describing visible objects near its current position. At the beginning of each episode, the agent is given a natural language instruction $I$ (e.g., Place the mug on the table and the book on the shelf). The instruction is grounded into a symbolic representation of the ground-truth goal $g$, denoted $g=\{(o_{i},l_{i})\}_{i=1}^{k}$, where each pair $(o_{i},l_{i})$ are the target object $o_{i}$ (e.g., mug) that should be placed at the specified location $l_{i}$ (e.g., on the table). In order to execute the instruction, the agent must internally derive the goal representation $\phi(I)\longrightarrow g$, where $\phi$ denotes the instruction grounding function, to guide the policy’s decision-making. Given an instruction $I$, where the policy $\pi$ is implemented by an LLM, the agent generates actions at timestep $t$ based on the trajectory of observations and actions:
$\pi(I,\tau_{t})\rightarrow a_{t},\quad\tau_{t}=(w_{1},a_{1},w_{2},a_{2},\ldots,w_{t-1},a_{t-1},w_{t})$ (1)
@@
The goal is to produce a sequence of actions $a_{1:t}=(a_{1},a_{2},\ldots,a_{t})$ such that the resulting state $s_{t}$ satisfies the agent’s goal $g(s_{t})=1$.
4 Memento
In this section, we introduce Memento, a personalized embodied agent evaluation framework designed to assess how well embodied agents leverage episodic memory containing personalized knowledge to provide personalized assistance. We begin by describing how we design the memory evaluation process (§ 4.1), categorize personalized knowledge (§ 4.2), the data construction process (§ 4.3), and the validation of our evaluation framework (§ 4.4).
Figure 2: Overview of Memento.
4.1 Memory Evaluation Process Design
Two-stage evaluation process.
The major challenge of evaluating embodied agents’ memory utilization capability is to quantify the memory effect on the overall task performance. To address this limitation, as shown in Figure 2, we divide the evaluation process into two stages.
  • Memory Acquisition Stage: Agents perform tasks with conventional object rearrangement task instructions containing personalized knowledge while accumulating the interaction history (i.e., episodic memory). The goal of this stage is to provide a reference performance baseline.
  • Memory Utilization Stage: Agents execute the same tasks as in the memory acquisition stage but with modified instructions that require agents to recall and apply the previously acquired personalized knowledge to succeed. The goal of this stage is to evaluate how well agents can utilize memory by comparing the performance drop relative to the acquisition stage.
Given the base object rearrangement task episode defined as the tuple $\epsilon=(S,I,g)$, the key concept of our evaluation process is to share the scene $S$ and goal representation $g$, while varying the instruction $I$ across the two stages to isolate instruction interpretation capability as the primary factor influencing performance differences. Formally, in the memory acquisition stage, each episode is defined as $\epsilon_{acq}=(S,I_{acq},g)$, where the instruction $I_{acq}$ contains sufficient information to infer the goal $g$, denoted as:
$\phi(I_{acq})\longrightarrow g$ (2)
During this stage, we also store the episodic memory $h_{acq}$:
$h_{acq}=(I_{acq},\tau_{k})\in H_{acq},\quad\tau_{k}=(w_{1},a_{1},w_{2},a_{2},\ldots,w_{k-1},a_{k-1},w_{k})$ (3)
In the subsequent memory utilization stage, each episode is defined as $\epsilon_{util}=(S,I_{util},g)$, where the instruction $I_{util}$ is intentionally underspecified and requires the agent to recall the corresponding episodic memory $h_{acq}$ to correctly interpret the goal $g$, formally:
$\phi(I_{util},h_{acq})\longrightarrow g$ (4)
Through this design concept, we are able to quantify the agent’s ability to utilize memory by comparing performance between the two stages.
Memory utilization stage task design.
Within our two-stage evaluation process design, we can assess embodied agents’ ability to utilize personalized knowledge from a single episodic memory. However, this approach alone fails to capture real-world complexity and lacks analytical diversity for comprehensive assessment. Therefore, to evaluate different levels of memory complexity, we divide our assessment into (1) Single-memory task, which require utilizing information from one episodic memory, and (2) Joint-memory task, which necessitate synthesizing information from two distinct episodic memories to successfully complete the episode. Building on our evaluation process concept, we form the joint-memory task by concatenating two episodes, which can be formulated as:
$\epsilon_{util}^{joint}=(S,I_{util}^{joint},[g^{i};g^{j}])$ (5)
where $i$, $j$ denotes the corresponding reference episodes from the memory acquisition stage.
4.2 Personalized Knowledge Categorization
Building upon our memory evaluation process, we aim to analyze embodied agents’ ability to understand personalized knowledge in object rearrangement tasks by focusing on its role in goal interpretation, $\phi(I)\rightarrow g=\{(o_{i},l_{i})\}_{i=1}^{k}$, where each $o_{i}$ and $l_{i}$ denote the target object and location, respectively. To facilitate this analysis, we categorize personalized knowledge into two types, each comprising subcategories  [^1] A detailed explanation of the sub-categories of personalized knowledge can be found in Appendix C.1. that reflect how users naturally express preferences and routines in real-world interactions. Each type is designed to isolate a distinct reasoning challenge that the agent must resolve by utilizing episodic memory during the memory utilization stage.
@@ 4.2 Personalized Knowledge Categorization
  • Object semantics: Individual objects that the user assigns personal meaning, encompassing subcategories such as ownership (e.g., my cup), preference (e.g., my favorite running gear), past history (e.g., a graduation gift from my grandma), or grouped references (e.g., my childhood toy collections). This category tests whether the agent can identify the target object $o_{i}$ by recalling its personal meaning from prior interactions.
@@
  • User patterns: Sequences of actions that the user consistently performs, including personal routines (e.g., my remote work setup) and arrangement preferences (e.g., my cozy dinner atmosphere) across recurring contexts. This category evaluates the agent’s ability to reconstruct the complete goal $g$ by leveraging previously observed behavioral patterns across multiple objects and locations.
4.3 Dataset Construction Process
We constructed the dataset for Memento through a four-step process using the Habitat 3.0 simulator [38] as the environment, with a simulated Spot robot as the agent [5, 58]. Our custom dataset spans 12 scenes, comprising a total of 438 episodes distributed across stages. Notably, the memory acquisition stage and the single-memory task in the utilization stage have the same number of episodes, whereas the joint-memory task contains fewer episodes. The detailed dataset statistics and the explanation of the construct process is described in Appendix C.3, and Appendix C.4.
Step 1: Object rearrangement task collection.
We use the test set of the PartNR [7] as our foundation object rearrangement task data. Unlike simple pick-and-place tasks, PartNR episodes require completing multiple object–location pairs within a single instruction, which aligns with our memory evaluation process design.
Step 2: Scene augmentation with distractor objects.
In the original scenes, there was an issue where no objects of the same type as the target object were present. As a result, the agent could identify the target object without needing to understand personalized knowledge. To address this, we augmented the scenes by placing distractor objects of the same type near the target object. For example, if the target object is a “blue cup” on the table, we place a “red cup” next to it as a distractor.
Step 3: Task instruction generation.
We first generate personalized knowledge contextually tailored to the original task instruction using GPT-4o. With the generated personalized knowledge, we applied to both stage instruction curation. As illustrated in Figure 2, the memory acquisition stage instruction $I_{acq}$ is constructed by concatenating the base instruction, object visual captions (only for episodes of the object semantics type), and the generated personalized knowledge. This ensures that the goal can be inferred directly from the instruction. For the memory utilization stage, we prompt GPT-4o to generate a personalized instruction that implicitly reflects the personalized knowledge, based on the same base instruction. For joint-memory tasks, we concatenate two such personalized instructions sequentially.

### v2

# Embodied Agents Meet Personalization: Investigating Challenges and Solutions Through the Lens of Memory Utilization

@@ Abstract
LLM-powered embodied agents have shown success on conventional object-rearrangement tasks, but providing personalized assistance that leverages user-specific knowledge from past interactions presents new challenges. We investigate these challenges through the lens of agents’ memory utilization along two critical dimensions: object semantics (identifying objects based on personal meaning) and user patterns (recalling sequences from behavioral routines). To assess these capabilities, we construct Memento, an end-to-end two-stage evaluation framework comprising single-memory and joint-memory tasks. Our experiments reveal that current agents can recall simple object semantics but struggle to apply sequential user patterns to planning. Through in-depth analysis, we identify two critical bottlenecks: information overload and coordination failures when handling multiple memories. Based on these findings, we explore memory architectural approaches to address these challenges. Given our observation that episodic memory provides both personalized knowledge and in-context learning benefits, we design a hierarchical knowledge graph-based user-profile memory module that separately manages personalized knowledge, achieving substantial improvements on both single and joint-memory tasks. Project website: https://connoriginal.github.io/MEMENTO
@@ 1 Introduction
Embodied agents empowered by large language models (LLMs) have recently demonstrated remarkable success in executing object rearrangement tasks in household environments (Huang et al., 2022b; Wu et al., 2023; Song et al., 2023; Chang et al., 2024; Li et al., 2025b). As the primary objective of embodied agents is to provide assistance to users while interacting with the physical world, leveraging LLMs’ natural language understanding and planning capabilities leads embodied agents to effectively interpret user instructions to successfully accomplish the task.
@@
But do such tasks truly reflect the challenges in providing practical assistance to the users? As illustrated in Figure 1, conventional embodied tasks predominantly focus on single-turn interactions with static and simplified instructions that the agents could simply follow without implicit reasoning to comprehend user intentions (Ahn et al., 2022; Huang et al., 2022a; Song et al., 2023; Wu et al., 2024). However, for personalized embodied agents, it is important to understand personalized knowledge that users assign unique semantics to the physical world (e.g., favorite cup, breakfast routine) to interpret dynamic instructions. To provide personalized assistance, agents must understand user instructions and conduct task planning by leveraging personalized knowledge retained from previous interactions—especially episodic memory, which enables the recall of specific events grounded in time and space (Huet et al., 2025). Therefore, understanding embodied agents’ memory utilization capabilities is crucial for providing practical assistance to users.
In this work, we investigate the challenges faced by LLM-powered embodied agents in personalized assistance tasks through the lens of memory utilization. We analyze agents’ memory utilization capabilities along two critical dimensions based on personalized knowledge: (1) the ability to identify target objects based on personal meaning (object semantics), and (2) the ability to recall sequences of object-location pairs from consistent user behavioral patterns, such as daily routines (user patterns). To evaluate task performance and quantify how memory utilization affects performance on personalized assistance tasks, we construct Memento, an end-to-end two-stage personalized embodied agent evaluation framework comprising both single-memory and joint-memory tasks. Using Memento, we address the following three research questions:
RQ1: Can current LLM-powered embodied agents effectively utilize memory to perform personalized assistance tasks?(Section 4) Initially, we evaluate various LLM-powered embodied agents using Memento to assess their memory utilization capabilities. Our findings reveal a key limitation: while agents can effectively recall simple object-related memory for object semantics, they struggle to apply sequential information such as user patterns to their planning process. Qualitative analysis shows that conflicts between parametric commonsense knowledge and non-parametric personalized knowledge can cause agents to neglect memory utilization, indicating that current LLM-powered embodied agents cannot reliably integrate personalized knowledge for task execution.
RQ2: What are the key factors that negatively affect embodied agents’ memory utilization capabilities?(Section 5) To understand the underlying factors that limit memory utilization performance, we conduct an in-depth analysis of agents’ behavior under varying memory conditions. Our investigation reveals two critical bottlenecks: (1) information overload: increasing the number of retrieved memories (top-$k$) introduces noise that degrades performance, and (2) coordination failures: agents cannot effectively coordinate multiple memories, failing even on simple joint-memory tasks combining two object semantic memories. These findings underscore the essential need for improved memory architectures that support agents in personalized assistance tasks.
@@
Figure 1: Comparison between conventional embodied tasks and personalized assistance tasks. Previous works focus on following simple instructions, while personalized assistance agents must know user-specific knowledge, which require grounding in past interactions.
RQ3: How can we design memory architectures to better support personalized assistance tasks?(Section 6) Based on our findings, we explore architectural approaches to improve agents’ memory utilization capabilities. Initially, we investigate memory summarization to remove unnecessary information but find marginal improvements and even degradation in smaller LLMs. This result reveals that episodic memory provides both personalized knowledge and in-context learning benefits, making it an essential memory module for embodied agents. Therefore, we develop a hierarchical knowledge graph-based user-profile memory module that manages personalized information independently. Adapting this memory module shows substantial improvements on both single and joint-memory tasks demonstrating the potential for future research in developing personalized embodied agents.
@@ 2 Related Work
LLM-powered embodied agents. LLMs have significantly advanced embodied agents’ reasoning and planning capabilities in recent years. Researchers have explored LLMs for interpreting user goals (Ahn et al., 2022), high-level task planning (Huang et al., 2022a), and integrating LLMs into comprehensive embodied agent frameworks (Huang et al., 2022b; Li et al., 2022; Mu et al., 2023; Song et al., 2023; Huang et al., 2023b). Other research directions have focused on generating executable code for embodied tasks directly from language instructions (Liang et al., 2022; Wang et al., 2023a; Singh et al., 2023), while various benchmarks have been developed to evaluate embodied reasoning abilities (Li et al., 2023; Choi et al., 2024; Chang et al., 2024; Li et al., 2025b). Collectively, these studies highlight the promise of LLM-powered agents in bridging language understanding and physical interaction.
@@
Memory systems for embodied agents. Previous studies on memory systems for embodied agents have primarily focused on semantic memory (e.g., scene graph, semantic map), which store and provide state information about the current environment (Rana et al., 2023; Kim et al., 2023; Gu et al., 2024; Xie et al., 2024; Wang et al., 2024c), or on procedural memory (e.g., skill library) that stores action primitives, focusing on how to perform tasks to enhance the efficiency in generating the low-level action code (Wang et al., 2023a; Sarch et al., 2023; Zhang et al., 2023). Another important category is episodic memory, which captures specific past interactions and experiences with users. However, prior uses of episodic memory have mostly treated it as passive task buffers (Ahn et al., 2022; Singh et al., 2023) or histories for in-context (Huang et al., 2022b; Liu et al., 2023; Song et al., 2023; Chen et al., 2024), without explicitly evaluating its role in personalized task grounding or systematic memory utilization.
@@
Personalization for embodied agents. The importance of personalization in robotics has long been recognized (Dautenhahn, 2004; Lee et al., 2012; Clabaugh & Matarić, 2018), particularly in the context of human-robot interaction where robots adapt their interactive behaviors to align with individual users. Recent works have focused on reflecting individual user’s preferences during embodied agents’ task execution, such as spatial arrangement (Kapelyukh & Johns, 2021; Wu et al., 2023), table settings (Puig et al., 2021), or personalized object navigations (Dai et al., 2024; Barsellotti et al., 2024). Recently, Xu et al. (2024) aim to infer user preferences from a few demonstrations and adapt planning behavior accordingly. However, these approaches primarily address implicit preference adaptation or short-term reactive behaviors derived from repeated user demonstrations, without leveraging explicitly provided, user-defined knowledge in a structured manner.
3 Evaluation Setup
3.1 Task & Focus
@@
Task: personalized object rearrangement. We formulate the personalized object rearrangement task as a Partially Observable Markov Decision Process (POMDP) $(S,A,T,R,\Omega,O,\gamma)$, where $S$ denotes the set of environment states, $A$ is the set of actions, $T$ is the transition function, $R$ is the reward function, $\Omega$ is the observation space, $O$ is the observation function, and $\gamma$ the discount factor. At timestep $t$, the agent receives a text observation $w_{t}\in\Omega$ describing nearby objects and generates actions via LLM policy $\pi$. Given a natural language instruction $I$ that may contain personalized references (e.g., "Place my favorite mug on the table"), the agent must derive the goal representation $g$ using both the instruction and personalized knowledge from memory: $\phi(I,M)\longrightarrow g={(o_{i},l_{i})}_{i=1}^{k}$, where $M={h_{1},h_{2},\ldots,h_{n}}$ represents the set of episodic memories from previous interactions, and each $(o_{i},l_{i})$ pair specifies target object $o_{i}$ and location $l_{i}$. The policy generates actions based on the observation-action trajectory $\tau_{t}$:
$\pi(I,M,\tau_{t})\rightarrow a_{t},\quad\tau_{t}=(w_{1},a_{1},w_{2},a_{2},\ldots,w_{t-1},a_{t-1},w_{t})$ (1)
@@
The goal is to produce action sequence $a_{1:t}$ such that the resulting state $s_{t}$ satisfies all object-location specifications in $g$.
Focus: memory utilization. Personalized assistance tasks require agents to integrate multiple capabilities including natural language understanding, memory retrieval, planning, and action execution. Among these interconnected abilities, we focus specifically on memory utilization—the agent’s capacity to effectively recall and apply relevant personalized knowledge from episodic memory $M$ to interpret instructions and guide task planning. This focus is motivated by memory utilization being the primary distinguishing capability from an agent perspective that separates personalized assistance tasks from conventional embodied tasks, making it essential to investigate for potential challenges and solutions for developing personalized embodied agents.
Figure 2: Overview of Memento. The framework evaluates memory utilization capability by comparing agent performance on tasks with identical goals but varying instructions on each stage.
3.2 Dimensions of Personalized Knowledge Analysis
We analyze embodied agents’ memory utilization capabilities along two dimensions that reflect how personalized knowledge affects goal interpretation in $\phi(I,M)\rightarrow g={(o_{i},l_{i})}_{i=1}^{k}$. Therefore, we categorize personalized knowledge into two types, each designed to isolate distinct reasoning challenges that agents must resolve using episodic memory:
@@ 3.2 Dimensions of Personalized Knowledge Analysis
  • Object semantics evaluates agents’ ability to identify target objects $o_{i}$ based on personal meaning assigned by users, including ownership (my cup), preference (my favorite running gear), personal history (a graduation gift from my grandma), or grouped references (my childhood toy collections). This dimension tests whether agents can recall specific object semantics from prior interactions to resolve ambiguous object references.
@@
  • User patterns assesses agents’ capacity to reconstruct complete goals $g$ by leveraging consistent behavioral patterns, including personal routines (my remote work setup) and arrangement preferences (my cozy dinner atmosphere). This dimension evaluates whether agents can apply previously observed patterns across multiple objects and locations to infer task requirements.
Each dimension targets different aspects of the grounding function $\phi(I,M)\rightarrow g$: object semantics focuses on resolving individual object references, while user patterns addresses holistic goal reconstruction from behavioral context. Detailed explanation of the sub-categories of personalized knowledge can be found in Appendix C.2.
3.3 Memento: Personalized Embodied Agent Evaluation Framework
Framework design. The major challenge of evaluating embodied agents’ memory utilization capability is quantifying the memory effect on overall task performance. To address this, as shown in Figure 2, we design an end-to-end two-stage evaluation process that measures memory utilization capabilities through instruction interpretation. Given the base object rearrangement task episode defined as the tuple $\epsilon=(S,I,g)$, the key concept of our evaluation process is to share the scene $S$ and goal representation $g$, while varying the instruction $I$ across the two stages to isolate instruction interpretation capability as the primary factor influencing performance differences.
In the memory acquisition stage, agents perform conventional object rearrangement tasks with instructions that contain sufficient personalized knowledge. These episodes are defined as $\epsilon_{acq}=(S,I_{acq},g)$ where $\phi(I_{acq})\longrightarrow g$, meaning agents can directly infer the goal from instructions. During this stage, agents establish a reference performance baseline while accumulating episodic memory $h_{acq}$ from their interactions. The memory utilization stage presents agents with identical tasks but uses underspecified instructions $\epsilon_{util}=(S,I_{util},g)$. These tasks can only be completed when agents recall and apply relevant personalized knowledge, requiring the extended grounding function $\phi(I_{util},h_{acq})\longrightarrow g$. Through this evaluation process, we are able to quantify the agent’s memory utilization capability by comparing performance between the two stages.
Furthermore, to evaluate different complexity levels, we assess both (1) single-memory tasks requiring information from one episodic memory, and (2) joint-memory tasks necessitating synthesis from two distinct memories, formulated as $\epsilon_{util}^{joint}=(S,I_{util}^{joint},[g^{i};g^{j}])$ where $i,j$ denote corresponding reference episodes from the acquisition stage. As shown in Figure 3, we conduct experiments without memory usage to validate our framework design, confirming that agents struggle significantly with underspecified instructions.